OpenStudy (anonymous):

If the cos(x) = -12/13, and tan(x) < 0, find sin(x) tan(x) sec(x

jhonyy9 (jhonyy9):

tan(x)=sin(x)/cos(x) so than sin(x)=tan(x)*cos(x) so than result that because cos(x) is negativ and tan(x) is less than zero than result that sin(x) will be always positive

OpenStudy (anonymous):

there is a right angled triangle with sides 5,12,13 and cos x = -12/13 and tan x < 0 means that x is in the second quadrant so sinx = 5/13 tan x = -5/12 secx = -12/5

OpenStudy (anonymous):

dont know why this last post was duplicated!